# MegaRAG: Multimodal Knowledge Graph-Based Retrieval Augmented Generation

### v1

@@ 4.1 Datasets
Global QA. To evaluate the global (book-level) QA capabilities of MegaRAG, we use two document collections: a textual corpus and a multimodal dataset. For the textual benchmark, we adopt the Ultradomain qian2024memorag dataset, which contains 428 college-level textbooks across 18 disciplines; we focus on four representative subsets: Agriculture (2,017,886 tokens), Legal (5,081,069 tokens), Computer Science (2,306,535 tokens) and Mixed-Domain (619,009 tokens). Since no standard benchmark exists for multimodal global QA, we curate a new multimodal benchmark comprising four documents: World History (a world history textbook, 788 pages), Environmental Report (a corporate environmental report slide deck, 422 pages), DLCV (an English lecture slide deck, 1,984 pages), and GenAI (a Chinese lecture slide deck, 594 pages).
@@
As these datasets lack manually labeled global questions, we adopt the question generation strategy from GraphRAG edge2024local and LightRAG guo2024lightrag. For each dataset, we use the document outline as input and prompt an LLM to create five synthetic RAG users, each with a profile describing their background and information needs. Each user is assigned five tasks representing distinct information-seeking goals, and each task is used to generate five questions that require a comprehensive understanding of the full document. This process yields 125 global questions per dataset.
Local QA. To evaluate local (slide- or page-level) QA, we use two benchmarks: SlideVQA yang2018hotpotqa and RealMMBench wasserman2025real. SlideVQA includes over 52,000 slides and 14,500 questions covering complex reasoning and numerical understanding, but its scale makes full evaluation computationally expensive. Instead, we construct a subset of 2,000 slides, referred to as SlideVQA (2k). RealMMBench assesses retrieval in multimodal RAG settings using visual-rich, table-heavy, and rephrased queries. RealMMBench consists of four sub-datasets: FinReport (2,687 pages), FinSlides (2,280 pages), TechReport (1,674 pages), and TechSlides (1,963 pages). Additional details are provided in Appendix A.
@@ 4.2 Baselines and Evaluation Metrics
As our approach is the first one automatically building Multimodal KGs for MMRAG-based question answering, we compare it with several widely adopted RAG baselines, including raw-source-based NaiveRAG, as well as KG-aided methods GraphRAG edge2024local, and LightRAG guo2024lightrag that are recent advancements in graph-based RAG. Details of them are provided in Appendix C. For fairness, besides the multimodal benchmark, we compare our method with them using only the textual benchmark too.
Global QA. In the absence of ground truth answers for global (book-level) questions, we follow the LLM-based evaluation strategy from GraphRAG edge2024local and LightRAG guo2024lightrag. Model responses are assessed along four qualitative dimensions: Comprehensiveness, Diversity, Empowerment, and Overall, as defined in prior work guo2024lightrag. Each response is compared against a baseline in a pairwise setup, with win rates (including ties) reported. Comprehensiveness measures how well the answer covers all aspects of the question; Diversity captures the richness and variety of perspectives; Empowerment reflects how effectively the answer informs and supports user understanding; Overall provides an aggregate score across the three preceding criteria.
Local QA. For local (slide- or page-level) QA, we evaluate performance by comparing the generated answers against ground truth answers. Specifically, LLM is used to judge whether the generated answer aligns semantically with the reference answer. Accuracy is then computed based on the proportion of correct matches. Further details regarding the evaluation dimensions and procedures are provided in Appendix C.
@@
A1 MegaRAG Tie A1 MegaRAG Tie A1 MegaRAG Tie A1 MegaRAG Tie
@@
A2 MegaRAG Tie A2 MegaRAG Tie A2 MegaRAG Tie A2 MegaRAG Tie
Comprehensiveness 0.0 100.0 0.0 0.8 88.0 11.2 0.0 98.4 1.6 0.0 100.0 0.0
Diversity 0.0 100.0 0.0 0.8 96.0 3.2 0.0 100.0 0.0 0.0 99.2 0.8
Empowerment 0.0 100.0 0.0 1.6 86.4 12.0 0.0 98.4 1.6 0.0 94.4 5.6
Overall 0.0 100.0 0.0 0.8 91.2 8.0 0.0 99.2 0.8 0.0 100.0 0.0
A3 MegaRAG Tie A3 MegaRAG Tie A3 MegaRAG Tie A3 MegaRAG Tie
Comprehensiveness 0.8 52.8 46.4 0.0 67.2 32.8 1.6 58.1 40.3 0.0 96.8 3.2
Diversity 12.0 72.8 15.2 8.0 79.2 12.8 7.3 80.6 12.1 0.8 97.6 1.6
Empowerment 5.6 70.4 24.0 4.0 77.6 18.4 16.9 57.3 25.8 0.8 99.2 0.0
Overall 1.6 61.6 36.8 0.8 75.2 24.0 3.2 70.2 26.6 0.0 98.4 1.6
Table 4: Ablation studies on four multimodal datasets in terms of win rates (%). A1: text-only graph construction (no visual inputs); A2: disable MMKG retrieval (page retrieval only); A3: replace two-stage generation with single-pass generation.
4.3 Implementation Details
To ensure consistency across all RAG methods, we standardize the LLM/MLLM implementation. Response generation and global question generation use GPT-4o-mini, while evaluation uses GPT-4.1-mini for greater robustness. All methods, including NaiveRAG, GraphRAG, and LightRAG, use OpenAI’s text-embedding-3-small model for textual embeddings. Textual documents are segmented into 1,200 token chunks with a 100-token overlap. We follow GraphRAG and LightRAG by setting their gleaning parameter to 1. The generation temperature is fixed at 0 across all tasks to reduce output variance.
For multimodal documents, we use the MinerU toolkit wang2024mineru to extract text, figures, and tables. MinerU converts PDFs into machine-readable formats while preserving layout and symbols, making it especially effective for processing scientific and technical documents. In MegaRAG, multimodal embeddings are encoded using GME-Qwen2-VL-2B zhang2024gme, which is designed to support a unified embedding space across single-, cross-, and fused-modality retrieval tasks. This allows MegaRAG to flexibly retrieve diverse input types within a consistent representation space. During retrieval, we set the top-$k$ value to $k=60$ for graph retrieval steps, following the dual-level retrieval strategy and set the top-$m$ value to $m=6$ for the page retrieval described in Section 3.2. For baselines without multimodal support, we retain only the extracted text and process it using the same pipeline as for textual documents. To mitigate inconsistencies, we standardize response prompts across all baselines, so output quality differences stem from model capabilities rather than prompt variations.
4.4 Main Results
Multimodal Global QA. An main characteristic of our method is that it can build MMKGs for RAG. In this experiment, we evaluate our MegaRAG on global QA tasks over multimodal documents. As shown in Table 2, MegaRAG outperforms all baselines on four visually rich datasets: World History, Environmental Report, DLCV, and GenAI. It achieves average win rates of 83.3% for Comprehensiveness, 92.7% for Diversity, 84.7% for Empowerment, and 89.5% Overall. The advantage is particularly evident on slide-based datasets such as DLCV and GenAI, where much of the core content is visual rather than textual. Compared with NaiveRAG and LightRAG, relying primarily on text, MegaRAG delivers stronger results across all evaluation dimensions. These gains stem from MegaRAG’s ability to build KGs that jointly encode textual information and visual cues.
@@ 4.4 Main Results
Multimodal Local QA. Table 3 shows the accuracy results on SlideVQA (2k) and the four RealMMBench subsets. Across all five test sets, MegaRAG performs more favorably. On SlideVQA (2k), which focuses on fine-grained slide-level reasoning, MegaRAG achieves 64.85% accuracy, higher than double the score of the strongest baseline. Similar trends are observed in RealMMBench. On FinSlides and TechSlides, which feature highly visual and table slide content, MegaRAG achieves 58.37% and 60.86%, outperforming the best baseline by 45 and 29 percents, respectively. Even in the more text-heavy FinReport and TechReport subsets, MegaRAG maintains a clear lead with 39.51% and 51.51%, surpassing LightRAG by 8 to 9 percents.
@@ 4.5 Ablation Study
To evaluate the contribution of each major component in MegaRAG, we conduct an ablation study by disabling key modules across the three main stages: MMKG construction, retrieval, and answer generation. In the first setting (A1), we remove all visual inputs, such as figures, tables, and page images, from the graph construction stage, relying solely on textual content. In the second setting (A2), we disable the MMKG-based retrieval mechanism and rely solely on the page retrieval. In the third setting (A3), we replace the two-stage generation pipeline with a single-pass generation setup that simultaneously considers both the subgraph and visual input.
(A1) Text-only graph construction. Removing visual inputs from the graph construction stage leads to a substantial performance decline across all datasets. Without visual entities and relations, the MMKG lacks critical cross-modal context, which is especially detrimental in visually rich domains such as GenAI. For example, the overall win rate on GenAI drops dramatically from 86.4% to just 0.8%. These results underscore the importance of incorporating visual elements in MMKG.
(A2) Disable MMKG retrieval. Disabling MMKG-based retrieval and relying solely on page retrieval results in the most severe performance degradation. Across all datasets and evaluation dimensions, MegaRAG achieves near 100% win rates when compared to this variant. This clearly demonstrates that structured retrieval over the MMKG is essential for accessing semantically rich and well-connected information, far outperforming page-level retrieval alone.
(A3) Remove two-stage answer generation. Replacing the two-stage generation pipeline with a single-pass setup causes moderate but consistent performance drops. Although this variant still benefits from MMKG construction and retrieval, average win rates decline by 14 to 25 percents. The largest drops appear in Diversity and Empowerment, suggesting that separating textual and visual reasoning before integration helps generate more nuanced and informative answers.
Among the three components, MMKG-based retrieval (A2) proves to be the most critical; its removal leads to a near-complete collapse in performance. Visual inputs in graph construction (A1) also play an important role, particularly for slide-centric documents, though their absence results in less dramatic losses. The two-stage generation strategy (A3) contributes more subtle but consistent gains, especially in generating diverse and empowering responses. Together, these results highlight the complementary value of all three components, with graph-based retrieval emerging as the core driver of MegaRAG’s effectiveness.
@@ A.1 Dataset Statistics
The Ultradomain benchmark (qian2024memorag) comprises 428 college-level textbooks spanning 18 academic disciplines. For this study, we focus on the four representative subsets:
@@
SlideVQA (2k). SlideVQA (tanaka2023slidevqa) includes over 52,000 slides and 14,500 questions covering complex reasoning and numerical understanding, but its scale makes full evaluation computationally expensive. Instead, we construct a subset of SlideVQA, which consists of 2,000 educational slides, featuring 1,581 figures, 139 tables, and 120,000 tokens.
The RealMMBench wasserman2025real is designed to evaluate retrieval performance in realistic multi-modal RAG scenarios, and contains four subsets:
@@ B.2 Retrieval and Generation Details
MegaRAG leverages the General Multimodal Embedder (GME) (zhang2024gme) to encode entities, relations, and page images within a unified embedding space. GME is built upon the Qwen2-VL architecture, a MLLM capable of processing text, images, or combined text–image inputs. It supports a broad range of retrieval tasks, including single-modality retrieval (e.g., text-to-text, image-to-image), cross-modality retrieval (e.g., text-to-image, image-to-text), and fused-modality retrieval (e.g., text with image to text with image). To generate embeddings, GME uses the final hidden state of the last token as the representation of the input. GME’s strength lies in its flexibility and generalization capability, making it well-suited for MegaRAG, which requires seamless integration of both text-to-text and text-to-page (image) retrieval tasks.

### v2

@@ 4.1 Datasets
Global QA. To evaluate the global (book-level) QA capabilities of MegaRAG, we use two document collections: a textual corpus and a multimodal dataset. For the textual benchmark, we adopt the Ultradomain Qian et al. (2024) dataset, which contains 428 college-level textbooks across 18 disciplines; we focus on four representative subsets: Agriculture (2,017,886 tokens), Legal (5,081,069 tokens), Computer Science (2,306,535 tokens) and Mixed-Domain (619,009 tokens). Since no standard benchmark exists for multimodal global QA, we curate a new multimodal benchmark comprising four documents: World History (a world history textbook, 788 pages), Environmental Report (a corporate environmental report slide deck, 422 pages), DLCV (an English lecture slide deck, 1,984 pages), and GenAI (a Chinese lecture slide deck, 594 pages).
@@
As these datasets lack manually labeled global questions, we adopt the question generation strategy from GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2025). For each dataset, we use the document outline as input and prompt an LLM to create five synthetic RAG users, each with a profile describing their background and information needs. Each user is assigned five tasks representing distinct information-seeking goals, and each task is used to generate five questions that require a comprehensive understanding of the full document. This process yields 125 global questions per dataset.
Local QA. To evaluate local (slide- or page-level) QA, we use two benchmarks: SlideVQA Yang et al. (2018) and RealMMBench Wasserman et al. (2025). SlideVQA includes over 52,000 slides and 14,500 questions covering complex reasoning and numerical understanding, but its scale makes full evaluation computationally expensive. Instead, we construct a subset of 2,000 slides, referred to as SlideVQA (2k). RealMMBench assesses retrieval in multimodal RAG settings using visual-rich, table-heavy, and rephrased queries. RealMMBench consists of four sub-datasets: FinReport (2,687 pages), FinSlides (2,280 pages), TechReport (1,674 pages), and TechSlides (1,963 pages). Additional details are provided in Appendix A.
@@ 4.2 Baselines and Evaluation Metrics
We compare MegaRAG with standard RAG baselines, including NaiveRAG and two recent KG-based methods, GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2025) (details in Appendix C). For fair comparison, we evaluate all methods on both the multimodal benchmark and the textual benchmark.
Global QA. Since global (book-level) questions lack gold answers, we adopt the pairwise LLM-judge protocol from GraphRAG/LightRAG Edge et al. (2024); Guo et al. (2025). Responses are evaluated by win rate (ties included) on four dimensions: Comprehensiveness, Diversity, Empowerment, and Overall Guo et al. (2025).
Local QA. For local (page/slide-level) QA with references, an LLM judges semantic equivalence between the generated and ground-truth answers, and we report accuracy. Additional evaluation details are provided in Appendix C.
4.3 Implementation Details
To ensure fair comparisons, we standardize the LLM/MLLM and prompting across all methods. Response generation and global question generation use GPT-4o-mini, and evaluation uses GPT-4.1-mini. All baselines (NaiveRAG, GraphRAG, LightRAG) use text-embedding-3-small for textual embeddings. Textual documents are chunked into 1,200 tokens with a 100-token overlap; for GraphRAG and LightRAG, we set gleaning to 1. We use temperature 0 for all tasks.
For multimodal documents, we use MinerU Wang et al. (2024a) to extract text, figures, and tables. MegaRAG encodes multimodal embeddings with GME-Qwen2-VL-2B Zhang et al. (2025). We retrieve top-$k$ entities/relations with $k=60$ for graph retrieval and top-$m$ pages with $m=6$ for page retrieval (Section 3.2). For text-only baselines, we keep only extracted text and apply the same textual pipeline.
4.4 Main Results
@@
VisRAG Ours Tie VisRAG Ours Tie VisRAG Ours Tie VisRAG Ours Tie
Comprehensiveness 0.8 35.2 64.0 4.0 31.2 64.8 4.0 27.2 68.8 0.0 87.2 12.8
Diversity 26.4 52.0 21.6 15.2 64.8 20.0 15.2 60.8 24.0 2.4 94.4 3.2
Empowerment 12.0 69.6 18.4 16.0 56.8 27.2 32.8 32.0 35.2 1.6 93.6 4.8
Overall 8.0 53.6 38.4 4.0 56.0 40.0 9.6 40.0 50.4 0.0 96.0 4.0
GME Ours Tie GME Ours Tie GME Ours Tie GME Ours Tie
@@
ColQwen Ours Tie ColQwen Ours Tie ColQwen Ours Tie ColQwen Ours Tie
Comprehensiveness 2.4 39.5 58.1 1.6 28.8 69.6 2.4 37.6 60.0 0.0 88.8 11.2
Diversity 29.0 57.3 13.7 14.4 65.6 20.0 17.6 57.6 24.8 1.6 94.4 4.0
Empowerment 14.5 62.9 22.6 16.0 53.6 30.4 25.6 44.8 29.6 1.6 90.4 8.0
Overall 6.5 50.8 42.7 3.2 52.0 44.8 9.6 50.4 40.0 0.0 94.4 5.6
Table 4: Pairwise LLM-judge win rates (%) of MegaRAG against MMRAG baselines on four multimodal datasets.
Multimodal Global QA. A main characteristic of our method is that it can build MMKGs for RAG. In this experiment, we evaluate our MegaRAG on global QA tasks over multimodal documents. As shown in Table 2, MegaRAG outperforms all baselines on four visually rich datasets: World History, Environmental Report, DLCV, and GenAI. It achieves average win rates of 83.3% for Comprehensiveness, 92.7% for Diversity, 84.7% for Empowerment, and 89.5% Overall. The advantage is particularly evident on slide-based datasets such as DLCV and GenAI, where much of the core content is visual rather than textual. Compared with NaiveRAG and LightRAG, relying primarily on text, MegaRAG delivers stronger results across all evaluation dimensions. These gains stem from MegaRAG’s ability to build KGs that jointly encode textual information and visual cues.
@@
Multimodal Local QA. Table 3 shows the accuracy results on SlideVQA (2k) and the four RealMMBench subsets. Across all five test sets, MegaRAG performs more favorably. On SlideVQA (2k), which focuses on fine-grained slide-level reasoning, MegaRAG achieves 64.85% accuracy, higher than double the score of the strongest baseline. Similar trends are observed in RealMMBench. On FinSlides and TechSlides, which feature highly visual and table slide content, MegaRAG achieves 58.37% and 60.86%, outperforming the best baseline by 45 and 29 percents, respectively. Even in the more text-heavy FinReport and TechReport subsets, MegaRAG leads with 39.51% and 51.51%, surpassing LightRAG by 8 to 9%.
4.5 Comparison with Multimodal RAG Methods
Table 4 presents pairwise LLM-judge win rates between MegaRAG and three strong multimodal RAG methods, namely VisRAG, GME, and ColQwen, on four multimodal documents.
MegaRAG consistently outperforms all three baselines across all datasets and evaluation dimensions, often by a substantial margin. While VisRAG, GME, and ColQwen mainly rely on multimodal retrieval over raw page images or unified multimodal embeddings, they do not explicitly model corpus-level structure or cross-page relations. By contrast, MegaRAG constructs multimodal knowledge graphs automatically, enabling structured abstraction and cross-page, multi-hop reasoning. This design yields particularly strong improvements in Comprehensiveness and Empowerment, where capturing dispersed evidence are important.
DLCV World History
A1 MegaRAG Tie A1 MegaRAG Tie
Comprehensiveness 6.4 49.6 44.0 0.0 72.0 28.0
Diversity 19.2 59.2 21.6 5.6 80.8 13.6
Empowerment 23.2 49.6 27.2 5.6 75.2 19.2
Overall 14.4 57.6 28.0 1.6 78.4 20.0
A2 MegaRAG Tie A2 MegaRAG Tie
Comprehensiveness 0.0 100.0 0.0 0.8 88.0 11.2
Diversity 0.0 100.0 0.0 0.8 96.0 3.2
Empowerment 0.0 100.0 0.0 1.6 86.4 12.0
Overall 0.0 100.0 0.0 0.8 91.2 8.0
A3 MegaRAG Tie A3 MegaRAG Tie
Comprehensiveness 0.8 52.8 46.4 0.0 67.2 32.8
Diversity 12.0 72.8 15.2 8.0 79.2 12.8
Empowerment 5.6 70.4 24.0 4.0 77.6 18.4
Overall 1.6 61.6 36.8 0.8 75.2 24.0
Table 5: Ablations on multimodal datasets (win rates, %). A1: text-only graph; A2: page-only retrieval; A3: single-pass (w/o two-stage) generation.
@@ 4.6 Ablation Study
We study the impact of MegaRAG’s core components by ablating modules in three stages: MMKG construction, retrieval, and answer generation. Table 5 reports the ablation results. (A1) removes all visual inputs (figures, tables, and page images) during graph construction, yielding a text-only graph. (A2) disables MMKG retrieval and uses only page retrieval. (A3) replaces our two-stage generation with a single-pass setup that jointly conditions on the retrieved subgraph and page images.
(A1) Text-only graph construction. Removing visual cues consistently degrades performance. Without visual entities and cross-modal links, the graph loses critical grounding signals, which is particularly harmful for visually intensive documents. This confirms the necessity of incorporating visual elements when building the MMKG.
(A2) Disable MMKG retrieval. This variant yields the largest drop. Relying solely on page retrieval substantially weakens evidence aggregation and multi-hop reasoning, and MegaRAG strongly outperforms this setting across evaluation dimensions. The results indicate that structured retrieval over the MMKG is the primary driver of our gains.
(A3) Remove two-stage answer generation. Using a single-pass generation causes moderate but consistent regressions. The largest decreases are typically observed in Diversity and Empowerment, suggesting that decoupling graph-based reasoning from visual grounding before fusion helps produce richer and more informative answers.
Overall, MMKG-based retrieval (A2) is the most critical component, while visual inputs in construction (A1) and the two-stage generation strategy (A3) provide complementary improvements.
Dataset (Pages) Input Tokens (Visual / Text) Metric GraphRAG (Total) MegaRAG
Init. Ref.
Total
(Init. + Ref.)
World History (788) 17.0M / 0.4M Time (min) 23.0 19.0 12.0 31.0
KG Tokens (M) 1.2 22.9 15.3 38.2
Env. eport (422) 25.0M / 0.2M Time (min) 10.0 8.5 5.5 14.0
KG Tokens (M) 0.6 32.6 21.7 54.3
Table 6: MMKG construction costs, where Init. and Ref. denote the Initial and Refinement stages.
4.7 Computational Overhead
We analyze the computational overhead of MegaRAG in terms of both indexing cost and inference latency.
Indexing Cost. MegaRAG constructs the MMKG in a one-time offline process consisting of an Initial Stage and a Refinement Stage. Table 6 reports the indexing costs on two representative datasets.
Importantly, graph construction is a one-time offline process; the resulting MMKG can be reused for subsequent queries without rebuilding the graph. Compared to text-only GraphRAG, MegaRAG incurs higher token usage and indexing latency (approximately $1.4\times$) due to processing image content (17M to 25M visual tokens). This additional cost enables MegaRAG to extract visual entities (e.g., 473 to 538 figures) that text-only methods cannot capture and to recover global dependencies missed by single-pass approaches.
Inference Latency. MegaRAG adopts a two-stage generation strategy. We measure end-to-end latency on the Environmental Report dataset using a single RTX 3090 (24GB); the latency can be further reduced with more powerful GPUs. MMKG retrieval and page retrieval run concurrently ($\sim$1.0s in total). In Stage 1, the MLLM generates answers from the retrieved subgraph ($\sim$20.5s) and from page images ($\sim$26s) in parallel. In Stage 2, a fusion step integrates both outputs ($\sim$16s), resulting in a total latency of $\sim$42s per question.
For comparison, GME, a strong MMRAG method, completes retrieval ($\sim$0.4s) and generation ($\sim$26s), resulting in a total latency of $\sim$26.4s. Although MegaRAG introduces an additional $\sim$15.6s latency, it produces MMKG-augmented answers that combine textual and visual evidence for stronger cross-page and multi-hop reasoning.
@@ 5 Conclusion
Limitations
Our experiments are primarily conducted on knowledge graphs constructed within a single book or report, and question answering across multiple books or reports remains limited. In addition, due to the high computational and preprocessing costs associated with figures and images, the scale of our multimodal dataset is still relatively limited.
Another limitation lies in our current treatment of visual content: each figure is represented as a single entity, which may overlook finer-grained elements or objects that are individually meaningful and worth representing as distinct entities. Moreover, since the MMKG is automatically constructed by MLLMs, hallucinations in entity or relation extraction may propagate to downstream reasoning. Understanding such hallucination behaviors and developing robust verification mechanisms are important directions for future work.
Acknowledgements
This work was supported in part by NSTC under Grants 112-2221-E-002-132-MY3 and 113-2634-F-002-003, NTU under Grant 114L900902, and the Esun-NTU collaboration project under Grant 114HZA3S002. We thank the National Center for High performance Computing (NCHC) of NARLabs in Taiwan for providing computational and storage resources.
Ethical Considerations
All data used in this work are publicly available documents. No human subjects or user-generated data are involved, and we do not collect or infer personally identifying information beyond what is already contained in the original sources. Our approach relies on LLMs, which may occasionally produce incorrect or biased text. In our setting, LLM outputs are primarily used to assist knowledge graph construction, while final answers are grounded in retrieved subgraphs and source pages and can be verified against the underlying documents, which helps limit potential errors.
Appendix
@@ A.1 Dataset Statistics
The Ultradomain benchmark (Qian et al., 2024) comprises 428 college-level textbooks spanning 18 academic disciplines. For this study, we focus on the four representative subsets:
@@
SlideVQA (2k). SlideVQA (Tanaka et al., 2023) includes over 52,000 slides and 14,500 questions covering complex reasoning and numerical understanding, but its scale makes full evaluation computationally expensive. Instead, we construct a subset of SlideVQA, which consists of 2,000 educational slides, featuring 1,581 figures, 139 tables, and 120,000 tokens.
The RealMMBench Wasserman et al. (2025) is designed to evaluate retrieval performance in realistic multi-modal RAG scenarios, and contains four subsets:
@@ B.2 Retrieval and Generation Details
MegaRAG leverages the General Multimodal Embedder (GME) (Zhang et al., 2025) to encode entities, relations, and page images within a unified embedding space. GME is built upon the Qwen2-VL architecture, a MLLM capable of processing text, images, or combined text–image inputs. It supports a broad range of retrieval tasks, including single-modality retrieval (e.g., text-to-text, image-to-image), cross-modality retrieval (e.g., text-to-image, image-to-text), and fused-modality retrieval (e.g., text with image to text with image). To generate embeddings, GME uses the final hidden state of the last token as the representation of the input. GME’s strength lies in its flexibility and generalization capability, making it well-suited for MegaRAG, which requires seamless integration of both text-to-text and text-to-page (image) retrieval tasks.